### v2

# CoTAL: Human-in-the-Loop Prompt Engineering, Chain-of-Thought Reasoning, and Active Learning for Generalizable Formative Assessment Scoring

Abstract
Large language models (LLMs) have created new opportunities to assist teachers and support student learning. Methods such as chain-of-thought (CoT) prompting enable LLMs to grade formative assessments in science, providing scores and relevant feedback to students. However, the extent to which these methods generalize across curricula in multiple domains (such as science, computing, and engineering) remains largely untested. In this paper, we introduce Chain-of-Thought Prompting + Active Learning (CoTAL), an LLM-based approach to formative assessment scoring that (1) leverages Evidence-Centered Design (ECD) principles to develop curriculum-aligned formative assessments and rubrics, (2) applies human-in-the-loop prompt engineering to automate response scoring, and (3) incorporates teacher and student feedback to iteratively refine assessment questions, grading rubrics, and LLM prompts for automated grading. Our findings demonstrate that CoTAL improves GPT-4’s scoring performance, achieving gains of up to 24.5% over a non-prompt-engineered baseline. Both teachers and students view CoTAL as effective in scoring and explaining student responses, each providing valuable refinements to enhance grading accuracy and explanation quality.
Index Terms:
Human-in-the-Loop, Formative Assessment, Automated Short Answer Scoring, Automated Grading, Prompt Engineering, LLM, LLMs, K12 STEM.
I Introduction
In K-12 STEM+C (Science, Technology, Engineering, Mathematics, and Computing) classrooms, educators foster engagement by linking scientific principles to real-world phenomena, enabling students to develop exploration, inquiry, and problem-solving skills. This provides students with opportunities to develop a foundational understanding of essential STEM+C concepts and practices [1]. Unlike single-discipline curricula, STEM+C learning requires students to integrate cross-domain concepts synergistically, linking ideas from one domain (e.g., science) to another (e.g., computing). While this approach enhances learning outcomes and promotes a deeper understanding of scientific processes, it also introduces complexities that can hinder learning [2], necessitating additional guidance and support that formative assessments can provide.
Assessments can incorporate open-ended questions that help students identify key learning constructs, apply learned concepts to problem-solving tasks, and develop critical thinking skills [3]. Simultaneously, they provide teachers with deeper insights into students’ STEM+C knowledge and problem-solving abilities [4]. Unlike summative assessments and standardized tests, which are primarily used for evaluation and may not capture the complexities students may face in achieving their learning goals [5], formative assessments are designed to support self-reflection when students have difficulties and facilitate feedback that helps them refine their understanding and improve their performance. At the same time, it allows teachers to monitor students’ learning progress and adjust their instruction to better meet student needs.
Evidence-based approaches for generating formative assessments and evaluation rubrics ensure alignment with curricular goals and designated standards [6]. Evidence-centered design (ECD) enables a more nuanced and flexible approach by adopting cognitive science and instructional design approaches that embed evidentiary reasoning into each stage of assessment and rubric development [5].
However, challenges persist in classroom environments for generating, grading, and providing timely feedback for formative assessments that align with educational standards. Designing and evaluating free-response formative assessments can place excessive demands on classroom teachers, who may have limited expertise in integrating STEM+C subjects, potentially affecting their ability to create assessments and rubrics that accurately evaluate students’ interdisciplinary knowledge [1]. Students may also lack mature writing skills, necessitating significant time and effort from human graders to infer the true implications of students’ answers [7]. Consequently, research is needed to develop automated scoring systems that efficiently deliver needed feedback to facilitate STEM+C learning based on teacher preferences [8].
Prompt engineering with large language models (LLMs) offers a promising solution, allowing users to adapt language models to downstream tasks by incorporating approaches like in-context learning (ICL) [9], chain-of-thought (CoT) prompting [10], and active learning [11] that obviate the need for traditional training via parameter updates, thus conserving computational resources. ICL enables LLMs to “learn” from labeled few-shot examples in the prompt during inference. CoT extends ICL by augmenting labeled few-shot examples with step-by-step reasoning chains to provide more explicit guidance to the LLM for scoring student answers [10]. Active learning is a human-in-the-loop approach to improving model training, where the human serves as an oracle, selecting additional instances to label for the next training iteration to enhance performance and robustness [11].
Unlike algorithmic prompt engineering, human-in-the-loop approaches [12, 13] integrate an important human perspective into the design and optimization pipeline that enable users to influence LLM outputs rather than relying solely on algorithmic decisions [14, 13]. This creates a human-AI collaboration [15] that enhances LLM alignment with user preferences and ensures the contextual relevance of the generated output by incorporating nuanced domain knowledge that algorithmic methods may overlook. Furthermore, this approach helps mitigate LLM errors and hallucinations by leveraging human expertise to validate and refine generated responses, fostering more accurate, ethical, and trustworthy LLM interactions.
In this paper, we pursue a stakeholder-AI partnership [^1] For the purposes of this paper, “stakeholders” refers to teachers, students, and researchers. to enrich assessment and feedback opportunities in a combined science, engineering, and computing NGSS-aligned curricular unit on Earth sciences [16, 17]. Building on prior work in automated scoring for short-answer science assessments, we introduce a generalizable human-in-the-loop LLM prompt engineering approach, Chain-of-Thought Prompting + Active Learning (CoTAL), to automate scoring and feedback for formative assessments across multiple domains. We define “generalizable” as the ability to accurately score formative short-answer responses that differ by question type (e.g., concept definitions, process descriptions, comparisons, and explanations), rubric structure (e.g., multi-label vs. multi-class), and content domain (science, computing, and engineering).
CoTAL consists of three phases: (1) leveraging ECD principles to design curriculum-aligned formative assessments and rubrics; (2) integrating these with human-in-the-loop prompt engineering via ICL, CoT, and active learning; and (3) refining the assessments, rubrics, and grading prompts through stakeholder feedback. This approach enhances LLM-based formative assessment response scoring and explanation capabilities while remaining grounded in principled assessment design.
Within this framework, we address the following research questions:
  • RQ1. Can CoTAL improve an LLM’s ability to score and explain responses to formative assessment questions across multiple connected domains?
  • RQ2. What do teacher and student input reveal about the effectiveness, actionability, and impact of CoTAL’s formative feedback?
We answer RQ1 using a mixed-methods approach. First, we conduct a quantitative evaluation of CoTAL’s scoring performance using GPT-4 on formative assessment questions in science, computation, and engineering, comparing it to a non-prompt-engineered baseline. We use Cohen’s Quadratic Weighted Kappa (QWK [18]) to measure agreement and assess CoTAL’s impact on scoring accuracy. Second, we qualitatively analyze GPT-4’s reasoning chains by performing a constant comparative analysis [19] to identify the strengths and weaknesses of the LLM’s scoring justifications when using CoTAL. We answer RQ2 qualitatively by conducting interviews with teachers and surveying students to assess the classroom effectiveness of LLM-generated formative feedback. We memo key findings from the teacher interviews [20] and apply constant comparative analysis to student survey responses, using stakeholder input to guide iterative methodological and curricular refinements.
Our results demonstrate that CoTAL generalizes effectively, significantly enhancing the performance of base GPT-4. Furthermore, CoTAL’s ability to clarify its scoring decisions builds trust in AI systems among students and teachers. This paper provides a pathway for introducing LLM-based assessment scoring in education and the learning sciences. The CoTAL approach can be generalized to show how LLMs can enhance automated formative assessment scoring and feedback generation while aligning with individual teacher preferences. For students, it provides additional learning opportunities by helping them reflect on their answers and overcome their difficulties.
II Background
LLMs offer new and exciting opportunities for addressing formative assessment grading and feedback generation [21, 22]. Previously, automated assessment scoring methods have incorporated data augmentation [23, 24], next sentence prediction [25], domain adaptation and supervised fine-tuning [26], prototypical neural networks [27], cross-prompt fine-tuning [28], and reinforcement learning [29] to improve grading accuracy. However, these methods have largely focused on more structured grading tasks in mathematics and computer science [30, 31], where open-ended responses are less prevalent than in science domains [12]. These approaches have achieved varying degrees of success but often fail to provide comprehensive insights into their scoring decisions.
Recently, researchers have used prompt engineering techniques to improve LLM performance on formative assessment scoring [12, 32]. However, in most cases, these approaches target a single domain and dataset, and their generalizability remains largely unevaluated. These approaches also fail to consider stakeholder input when developing and refining their methods. For example, Lee et al. (2024) [33] demonstrates the effectiveness of CoT prompting using contextual item stems and rubrics, emphasizing the importance of domain-specific reasoning in improving LLM performance when scoring middle school science assessments. Their prompt engineering procedure, WRVRT (Writing, Reviewing, Validating, Revising, and Testing), leverages CoT prompting and iterative refinement to enhance scoring accuracy and explainability. However, their study focuses solely on the science domain, and it remains unclear how well their approach generalizes to other subjects. Their work also centers on the prompt engineering process, without incorporating ECD principles or integrating feedback from students and teachers to refine their systems.
There exists a notable gap in automated formative assessment scoring approaches that (1) integrate ECD design principles with prompt engineering, (2) explain scoring decisions to teachers and students, (3) are effective across multiple domains, and (4) incorporate stakeholder feedback for refinement. In this paper, we address these gaps by proposing a generalizable, stakeholder-informed framework that combines ECD with human-in-the-loop prompt engineering to support transparent, accurate, and adaptable formative assessment scoring. Our work builds on prior research by explicitly grounding our methodology in assessment design theory, engaging teachers and students to refine our approach and evaluating performance across science, engineering, and computing domains.
III SPICE Curriculum
SPICE is a three-week middle school curriculum unit that challenges students to redesign their schoolyard using surface materials to minimize the amount of water runoff after a rainstorm while adhering to design constraints (such as considering construction cost and accessibility needs). The problem-based learning curriculum comprises five core units: (1) physical experiments; (2) conceptual modeling; (3) paper-based computational thinking tasks; (4) computational modeling of the water runoff phenomenon; and (5) engineering design, where students use their computational models developed in SPICE’s computer-based learning environment to redesign a schoolyard while considering engineering constraints [34]. The curriculum targets NGSS performance expectations for upper elementary and middle school Earth science and engineering design, emphasizing surface water movement and human impact.
Formative assessments are interspersed throughout the curriculum to help students self-evaluate their learning progress. These assessments help teachers monitor and support students’ science, computing, and engineering learning as they progress through the curriculum. The associated grading rubrics (discussed shortly) reflect cross-domain connectivity, enabling teachers to monitor students’ progress and modify instruction when students have difficulties.
SPICE leverages ECD to systematically create assessments and tasks to evaluate student learning in science, computing, and engineering. For our analysis, we picked three of the seven formative assessments (F2, F3, and F5) developed to monitor and support students’ learning across the three domains in the curriculum. The locations of the assessments in the integrated curriculum are shown at the identified markers in Figure 1.
Figure 1: SPICE curricular sequence (L items in white are lessons; F items in red are formative assessments.
Previous work focused on the automated grading of Science Concepts and Reasoning Task assessments from the science unit (F1). This paper generalizes the previous approach to include three additional formative assessments that cover:
  1. 1.
    a Rules Task (F2) that requires students to structure their understanding of conservation of matter by expressing the relation between the amounts of rainfall, water absorbed by the surface material, and runoff as three separate “rules”;
  2. 2.
    a Debugging Task (F3) requiring students to check the conditional form of the rules and value expressions from the Rules Task by analyzing block-based code generated by a fictional classmate and identifying errors in the computational model; and
  3. 3.
    an Engineering Task (F5) where students integrate their knowledge of science and computing concepts with the engineering principle of fair tests to ensure a fair comparison between two designs (i.e., the conditions under which the designs are evaluated remain consistent).
These assessments (and their rubrics) were explicitly designed for students to realize the connections between the science, computing, and engineering domains and to form a cumulative understanding of cross-domain conceptual knowledge as they progress through the curriculum [35, 17].
[…]
Subscore Description Domain
R1 Completed if statement “if rainfall is less than absorption limit.” SCI, COM
R2 Set absorption to rainfall in this rule. SCI
R3 Set runoff to 0 in this rule. SCI
R4 Completed if statement “if rainfall is equal to absorption limit.” SCI, COM
R5 Set absorption to rainfall OR absorption limit in this rule. SCI
R6 Set runoff to 0 in this rule. SCI
R7 Completed if statement “if rainfall is greater than absorption limit.” SCI, COM
R8 Set absorption to absorption limit in this rule. SCI
R9 Set runoff to “rainfall - absorption limit” OR “rainfall - absorption” in this rule. SCI
TABLE I: Categorical rubric used for the Rules Task. Each “R” corresponds to a different subscore for the Rules Task and is explained in the table. The Rules Task targets the science (SCI) and computing (COM) domains.
Table I enumerates the nine possible points (subscores) for the Rules Task (R1-R9). Students receive 1 point per correct conditional statement they identify (R1, R4, R7). Within each conditional statement, students receive 1 point for correctly setting the absorption value (R2, R5, R8) and 1 point for correctly setting the runoff value (R3, R6, R9). For example, the statement “if rainfall is equal to the absorption limit, then set absorption to rainfall, and set runoff to zero” would earn 3 points (R4, R5, and R6).
Figure 2: Erroneous code presented to students during the Debugging Task. Red “D” circles correspond to the individual model errors presented in Table II.
For F3, the Debugging Task, students are asked to identify and describe the five errors present in a fictional student’s code (illustrated in Figure 2). To accomplish this, students must understand the conditional logic expressions they wrote for the Rules Task and accurately translate them into block-structured programming code using “if statements” and expressions in the SPICE environment to model the amount of absorbed water and runoff. Students learn how to use the coding blocks to build their computational water runoff models before they work on the Debugging Task. The rubric for this task is shown in Table II. This paper uses the terms “model” or “computational model” to refer to the block-based code representation and “language model” or “LLM” to refer to GPT-4 (discussed in Section IV-B).
Subscore Description Domain
D1 “Set absorption limit” should be before the first conditional statement. COM
D2 In the “less than” condition, rainfall should be compared to the absorption limit. SCI, COM
D3 In the “less than” condition, absorption should be set to rainfall. SCI
D4 The “greater than” condition should not be embedded in the less than condition, but connected to it. COM
D5 In the “greater than” condition, absorption should be set to absorption limit, not the other way around. SCI, COM
TABLE II: Categorical rubric used for the Debugging Task. Each “D” corresponds to a different subscore (i.e., code error; see Figure 2) for the Debugging Task and is explained in the figure. Like the Rules Task, the Debugging Task targets the science (SCI) and computing (COM) domains.
Table II shows that students can earn up to 5 points (subscores; one for each error they identify) for the Debugging Task. “D” values refer to the individual errors the students must identify in the model. D1 refers to the “set absorption limit (inch) of the selected material” block being erroneously placed on line 6 (it should come before the first conditional statement on line 3). D2 refers to rainfall being incorrectly compared to absorption in the “less than” condition on line 7 (it should be compared to absorption limit). D3 refers to absorption incorrectly being set to the absorption limit inside the “less than” condition on line 8 (absorption should be set to rainfall). D4 refers to the “greater than” condition being improperly set on line 10 (it should not be nested inside the “less than” condition). D5 refers to the absorption and absorption limit being swapped inside the “greater than” condition on line 11 (absorption should be set to absorption limit, not the other way around).
The Engineering Task formative assessment (F5) requires students to integrate their science and computing domain knowledge with their engineering knowledge of design constraints and fair tests. Students compare two design solutions generated by a fictional student and are provided information about each design test’s input (e.g., rainfall) and output (e.g., cost and runoff). Students are then asked to explain whether the provided information allows them to conclude that one design is better than the other.
This task utilizes a new rubric structure where students are awarded a single numerical score from 0 to 4 points. Students are assessed on their ability to determine if the reported tests allow a valid comparison between the two design solutions. Since the fictional student uses different rainfall values to compare the runoff between the designs, the comparison is not “fair” because the outcome variable is not generated using the same input for both tests. The students’ explanations are evaluated at different levels using the rubric in Figure III.
Score Description Domain
4 Student explains that (1) the designs cannot be compared because different rainfall values were used to test each one, and (2) the runoff comparisons will not be “fair.” ENG, SCI, COM
3 Student explains the designs cannot be compared because different rainfall values were used to test each one. ENG, SCI
2 Student explains the designs cannot be compared because both tests violate design constraints, demonstrating an understanding of constraint satisfaction but not the need for fair tests. ENG
1 Student identifies that the designs cannot be compared but does not provide reasoning. ENG
0 Student answers “yes” that both designs can be compared fairly. SCI, COM
TABLE III: Engineering Task rubric, targeting the science (SCI), computing (COM), and engineering (ENG) domains.
Table III shows that students are awarded 0 points if they fail to recognize that the two designs cannot be compared fairly (i.e., they answer “yes” to the question posed). Students receive one point if they identify that the two tests cannot be compared fairly (i.e., they answer “no” to the question posed) but do not provide a meaningful explanation. Students receive two points if they discuss design constraints as a reason the two tests are not comparable. Three points are awarded if the students discuss the differing rainfall values as the reason why the tests cannot be compared fairly. Four points are awarded if the students mention the differing rainfall values and explain that this results in unfair runoff comparisons.
IV Methods
IV-A Chain-of-Thought Prompting + Active Learning (CoTAL)
We have developed CoTAL (illustrated in Figure 3), a generalizable method for improving automated formative assessment scoring and aligning LLM scores and explanations with the needs of teachers and students. Our approach fosters a collaborative partnership among researchers, teachers, students, and AI, integrating ECD, human-in-the-loop prompt engineering, and stakeholder feedback to (1) score and explain students’ short-answer responses to formative assessment questions in the SPICE curriculum; and (2) iteratively refine our formative assessments, rubrics, and grading prompts based on student and teacher feedback.
Figure 3: Chain-of-Thought Prompting + Active Learning (CoTAL).
CoTAL is structured into three distinct phases, as shown in Figure 3. In Phase I (yellow), teachers and researchers collaborate using ECD to design formative assessments and rubrics that align with curricular goals. The rubrics are developed to evaluate students’ responses to each formative assessment. Phase I is human-driven, with both parties (teachers and researchers) working collaboratively to create the formative assessments and their corresponding grading rubrics. The LLM is not considered in Phase I. This study’s Phase I formative assessments and rubrics are detailed in Section III.
In Phase II (blue), researchers work with the LLM to develop and optimize an initial prompt by (1) sampling a subset of student responses and conducting IRR to establish scoring consensus, (2) employing few-shot ICL and CoT reasoning to align the LLM with human consensus, and (3) iteratively refining the prompt through active learning by generating scores and explanations using a validation set to ensure the LLM’s generations align with human scoring preferences. These steps correspond to the blue diamonds, which represent Response Scoring, Prompt Development, and Active Learning in Figure 3. This procedure was applied to the Science Concepts and Reasoning Task in previous work [12], as well as the three new integrated assessments in science, computing, and engineering discussed in this paper (see Section III).
In Step 1, Response Scoring, human experts independently score a random subset of student responses using the rubrics developed in Phase I and establish their initial degree of agreement based on an IRR metric (e.g., Cohen’s Kappa [36]). When discrepancies arise between the scores assigned by the two reviewers, the underlying reasons for these differences are analyzed and resolved to reach a final consensus. Difficult-to-resolve discrepancies, termed “sticking points,” receive special attention during Prompt Development (Phase II, Step 2) to refine the prompt structure and guide the LLM toward alignment with the human scorers. Sticking points often emerge from “edge cases” where reviewers interpret the same rubric differently. For example, one Engineering Task sticking point was whether a general understanding of design constraints (i.e., the student did not explicitly mention any constraint by name) qualified for 2 points (both reviewers ultimately agreed this was acceptable). This process—randomly sampling the data, calculating Cohen’s $k$, and documenting reasons for disagreement—is repeated until $k\geq 0.70$ to ensure consensus. This is depicted as the “IRR” self-loop in Figure 3.
In Step 2, Prompt Development, the LLM is first given task instructions, followed by the formative assessment question, grading rubric, and additional context to guide the LLM’s scoring. This supplementary information helps the LLM connect student responses to the rubric and comprehend the intended concepts, questions, and scoring criteria. Additionally, various prompt patterns are employed to guide the LLM during inference and help structure its output, including a: (1) persona pattern that assigns the LLM a “persona,” or role, to play while generating its output; (2) context manager pattern that defines the context the LLM should consider while generating its output; (3) template pattern that provides the LLM with a structured output template; and (4) meta language creation pattern that creates a custom language for the LLM to understand (e.g., a textual shorthand notation to help the LLM interpret graphs) [37]. For example, the context manager pattern helped the LLM understand the relationship between the Rules and Debugging Tasks (i.e., the students often referred to conditional statements by “rule” number during the Debugging Task), and the meta language creation pattern allowed us to distill the Debugging Task’s computational model image into LLM-readable text.
After constructing the prompt, labeled few-shot instances, along with CoT reasoning chains, are added to it to align the LLM with the consensus reached by human scorers. Each formative assessment prompt features two types of few-shot examples: (1) ground truth instances, where both scorers agree during Response Scoring; and (2) sticking point instances, where scorers disagree during Response Scoring and the reasons for disagreement carry over to other instances, leading to similar disagreements among the scorers. For sticking point instances, CoT reasoning chains help clarify potential misunderstandings and guide the LLM toward human consensus. All few-shot instances are accompanied by CoT reasoning chains that loosely (i.e., not verbatim) follow this template:
The student says X. The rubric states Y. Based on the rubric, the student earned a score of Z.
Unlike traditional CoT prompting, which relies on the LLM to generate intermediate reasoning chains based solely on patterns learned during training, CoTAL grounds the LLM’s responses by instructing it to cite relevant portions of students’ answers verbatim and link them to the scoring criteria in the rubric. In this way, the LLM’s scoring decisions and explanations are inherently guided by human input during inference, ensuring generations remain faithful to the established criteria.
In addition to incorporating ground truth and sticking point examples in the prompt, additional instances are included to ensure that few-shot examples are proportionally represented across subscores to achieve data balance. Min et al. (2022) [38] recommend balancing few-shot instances based on the true distribution of the dataset’s labels rather than doing so uniformly. However, in multi-label datasets such as ours, adding individual instances can shift the label distribution for each subscore category, making perfect balance difficult to achieve. At a minimum, few-shot examples should include at least one positive and one negative instance for each subscore (multi-label) or one instance of each score (multi-class).
Once the prompt is constructed, Active Learning (Step 3) tests the prompt against a validation set. The LLM’s generations are analyzed by first isolating the incorrectly scored instances. For each subscore that the LLM predicts incorrectly, we identify a “scoring trend” to determine whether the LLM tends to produce false positives (FPs) or false negatives (FNs) and qualitatively discern the reasons behind these inaccuracies. This process mirrors Response Scoring, where recurring LLM errors on the validation set are identified as sticking points. We then select validation set instances exemplifying these sticking points, annotating them with CoT reasoning chains to correct the LLM’s mistakes and adding them to the existing few-shot examples in the prompt (illustrated by the “AL Prompt Refinements” loop in Figure 3).
Phase II of CoTAL shares similarities with explainable AI (XAI) approaches, which focus on explaining a model’s decisions based on its internal mechanisms. However, in CoTAL’s case, the explainability comes from the LLM’s generations, offering insights into how the model evaluates student responses in relation to rubrics. The emphasis is on alignment with grading criteria rather than interpreting the model’s internal logic. This explainability is particularly valued by students, who see it as a critical factor for building trust in AI systems in educational settings (see Section VI). Unlike our prior work, which primarily highlighted our human-in-the-loop prompt engineering contributions, this paper “closes the loop” (see Figure 3) by (1) integrating LLM prompt engineering with ECD principles, (2) investigating our method’s practical utility in classrooms through studies involving both teachers and students, and (3) leveraging student and teacher feedback to inform refinements to our formative assessments, rubrics, and prompts.
In this work, we refine our original approach based on prior findings [12]. Previously, we conducted active learning by inserting several validation instances back into the prompt. However, we found that the LLM tended to overfit when the number of few-shot instances was large, or the CoT reasoning chains became too granular. To mitigate the LLM’s tendency to overfit, we insert only a single instance into the prompt during Active Learning in this study, specifically targeting the most persistent LLM errors. Additionally, we use Prompt Development (Phase II, Step 2) to provide the LLM with a list of “guidelines” [^2] In the actual prompt, we refer to the guidelines as “rules,” but we use the term “guidelines” in the manuscript so as not to be confused with the “rules” in the Rules Task. that the LLM is instructed to adhere to at all times. For example, one Rules Task guideline instructed the LLM that the order in which students listed the three rules should not affect its scoring decisions. Human scorers agree upon these guidelines based on the consensus they reach during Response Scoring. Just as we use the CoT reasoning chains to cite evidence from the student’s response and tie that evidence to the rubric to elicit a score, we similarly use the CoT reasoning chains to cite these guidelines to improve LLM responses.
In Phase III, the prompt refined during Phase II is deployed in a classroom setting to score students’ formative assessment responses and explain the assigned scores. Researchers then sample the LLM’s responses and present them to teachers, who critique the model’s scoring accuracy, explanatory soundness, and clarity. The teachers and researchers identify the LLM’s strengths and weaknesses during these discussions. Based on this feedback, they collaborate to determine how to best address the LLM’s shortcomings without compromising its strengths, agreeing on specific refinements to the formative assessment questions, scoring rubrics, and prompts.
Similarly, students are shown the LLM’s responses to their formative assessment answers and are asked to critique them. While teacher feedback informs methodological improvements related to curricular goals, student feedback highlights user experience and personalized learning. For example, students might emphasize elements such as the LLM’s tone in its responses and the overall effectiveness of the content it provides to enhance their understanding of relevant concepts. This process is illustrated in Figure 3 by the “Curricular and Instructional Refinements” and “Teacher/Student Prompt Refinements” loops. Like Phase I, Phase III depends on human input, involving close collaboration among researchers, students, and teachers to guide curricular and methodological enhancements.
IV-B Experimental Design
We analyzed formative assessment responses from 175 6th-grade students (ages 11-12) at a public middle school in the southeastern United States. The student population was 67% White, 14% Black/African American, 11% Asian, 8% Hispanic/Latino, and included three students of other races, with a near-equal gender distribution of 51% male and 49% female. The data collection and analysis protocol was approved by Vanderbilt University’s IRB, and all students whose data was analyzed provided consent. While 175 students participated in the study over two years, several students’ formative assessment responses were omitted from our analysis at various points due to either non-consent or absences resulting in incomplete work. In total, 158 instances were available for the Rules Task, 166 instances were available for the Debugging Task, and 161 instances were available for the Engineering Task.
During Response Scoring, two of this paper’s authors independently scored a randomly selected subset (20%) of the data using the rubrics described in Section III. For each formative assessment, the two humans compared scores and discussed their disagreements before reaching a consensus. Cohen’s Kappa, the predominant measure in the literature assessing inter-rater reliability between two reviewers, was used as the IRR measure. The Cohen’s $\kappa$ values for the Rules, Debugging, and Engineering Tasks were 0.861, 0.740, 0.844, respectively.
Once consensus was achieved ($\kappa\geq 0.70$), one of this paper’s authors scored the remaining instances. Each task’s dataset was split into 80/20 training/testing sets before Prompt Development, with training set instances being those considered for inclusion in the prompt as few-shot examples and test set instances used for method evaluation. Instances discussed during IRR were withheld from the test set to prevent data leakage. Instances in the training set not used as few-shot examples in the initial prompt were reserved as a validation set to support the active learning process.
The datasets for all three tasks were imbalanced. In the Rules Task, the mode for total score was a perfect 9/9, occurring in 34 out of 158 cases, while a score of 0 was the second most frequent (observed in 29 instances). While examining the six subscores related to conditional statements and runoff values, a majority of students earned points. Conversely, for the three subscores related to absorption values, students were more likely not to receive points, suggesting students generally understood the logic expressions and runoff values but struggled to understand their relationship to absorption. The Debugging Task’s distribution was biased towards higher scores, with the mode being a perfect 5/5 for 66 out of 166 students and demonstrating a decreasing frequency of students attaining lower scores. All five Debugging Task subscores resulted in students earning points more often than not. The Engineering Task was predominantly characterized by incorrect responses, with the most common score being 0 (accounting for 63%; 101 out of 161). The remaining scores were evenly distributed across the range from 1 to 4 (inclusive).
To evaluate CoTAL’s scoring accuracy on the Rules, Debugging, and Engineering Tasks, we compared the performance of CoTAL to a zero-shot, “scoring-only” Baseline (i.e., numerical scores only without labeled instances, CoT reasoning chains, or active learning) to evaluate CoTAL’s generalizability, comparing CoTAL-generated scores and explanations to those of the non-prompt-engineered LLM. Performance details for adding each individual component to CoTAL’s prompt engineering pipeline are reported in previous work [12]. We used GPT-4333A temperature of 0 was used to achieve near-deterministic behavior. to conduct our analysis due to its balance between cost and accuracy.
To assess CoTAL’s generalizability for scoring and explaining formative assessment questions across multiple connected domains (RQ1), we adopted a mixed-methods approach. First, we measured LLM performance on a test set quantitatively, using Cohen’s QWK due to its prevalence in the automated essay scoring literature [39]. Unlike traditional Cohen’s $\kappa$, Cohen’s QWK awards partial credit based on the degree of disagreement, making it ideal for ordinal data. Next, we qualitatively investigated CoTAL’s impact on GPT-4’s scoring accuracy and its ability to provide useful feedback by conducting a constant comparative analysis of the LLM’s scoring explanations and errors to determine CoTAL’s strengths and weaknesses. CoTAL details for each of the three formative assessments are provided in our Supplementary Materials [^4] https://github.com/claytoncohn/TLT25_Supplementary_Materials, and we discuss our findings for RQ1 in Section V.
To examine student and teacher impressions of CoTAL’s feedback efficacy and impact in supporting classroom learning (RQ2), we employed qualitative analysis by first conducting two semi-structured interviews with two classroom teachers. Each teacher had over five years of experience teaching the SPICE curriculum and more than 20 years of overall teaching experience. Teachers reviewed LLM responses to previously unseen Science Concepts and Reasoning (F1 in Figure 1) assessment questions, expressing their agreement with the LLM’s scores and explanations and sharing their preferences regarding the response structure. Additionally, we asked teachers how we could improve LLM outputs to further benefit both students and teachers.
Second, we ran a separate focus group study with a subset of our study participants (23 students) to evaluate CoTAL’s performance in scoring their Science Concepts and Reasoning assessment. The students reviewed the AI-generated scores and feedback and then completed a survey. The survey assessed their agreement with GPT-4’s scoring accuracy, its feedback utility, and their confidence in the system’s ability to grade future assignments. To answer RQ2, we (1) created memos of key findings from the teachers’ interviews and (2) conducted a constant comparative analysis of the students’ survey responses. RQ2 results are presented in Section VI.
V Analyzing RQ1: CoTAL Generalizability Across Multiple Connected Domains
RQ1 asked, Can CoTAL improve an LLM’s ability to score and explain responses to formative assessment questions across multiple connected domains? To answer this question quantitatively, we first present performance comparisons between CoTAL and the Baseline for the Rules, Debugging, and Engineering tasks. We then present our qualitative findings, identifying CoTAL’s strengths and weaknesses for each task.
V-A Rules Task
Performance results for the Rules Task comparing CoTAL and the Baseline are presented in Table IV.
Rule Baseline CoTAL
R1 0.840 1.000
R2 0.936 1.000
R3 0.934 0.934
R4 0.467 0.784
R5 0.875 0.813
R6 1.000 1.000
R7 0.632 0.840
R8 0.934 0.934
R9 0.811 0.938
Total Score 0.930 0.968
TABLE IV: Performance results for CoTAL applied to the Rules Task relative to the Baseline in terms of Cohen’s QWK. Each “R” corresponds to a different subscore for the Rules Task and is explained in Table I. Total Score compares the LLM’s prediction of the total score (i.e., the sum of all 9 subscores) to the human label. For each metric, the best-performing implementation is in boldface.
The Rules Task Baseline implementation resulted in an average QWK of 0.826 across all subscores. Applying CoTAL resulted in an average QWK of 0.916, which represents an average increase of 0.090 (10.9% ) over the Baseline while using CoTAL. By itself, the Baseline implementation provided good results with 7/9 Baseline subscores showing “strong” agreement [40]. Incorrect predictions were evenly split between false positives and negatives (7 and 6, respectively). The Baseline underscored the Total Score by exactly one point in 7 of the 32 test set instances because it failed to properly add the predicted subscores (e.g., the LLM awarded 8 subscore points but a Total Score of 7). Interestingly, this “one under” issue was not observed in previous work [12]. We hypothesize that a large number of subscores (nine for the Rules Task) may cause issues during inference due to the exponential nature of combinatorics, resulting in $5122^{9}=$ 512 scoring combinations. Further research is necessary to evaluate this hypothesis empirically. The LLM Baseline incorrectly predicted 17/32 Total Scores overall (12 of which were underscores).
CoTAL improved results, with every individual subscore and Total Score (except R5) beating or tying the Baseline. CoTAL reduced false positives in individual subscores from 7 to 1 while only increasing false negatives by 1. This was explicitly addressed during Active Learning, which favored false positives to false negatives by a 2:1 ratio. R5 was the only subscore with more than one incorrect prediction with CoTAL, while the Baseline had 4/9 subscores with multiple incorrect predictions. CoTAL resulted in 10 incorrectly predicted Total Scores, compared to 17 for the Baseline implementation, with 8 instances of underscoring and 2 of overscoring. Only 2 of the 10 incorrectly predicted Total Scores were off by two or more points, compared to 5 in the Baseline approach. CoTAL also reduced the Baseline’s “one under” Total Score addition errors from 7 with the Baseline to 2 using CoTAL.
In addition to accurate scoring, the LLM provided accurate feedback for the Rules Task, citing relevant evidence, tying it to specific rubric components, and justifying its scoring decisions like a human scorer. The following example highlights a successful LLM generation for R4 in the Rules Task.
Student: If rainfall is greater than absorption limit set runoff to absorption limit minus rainfall, and set absorption to absorption limit.
GPT-4: The student says ‘set runoff to absorption limit minus rainfall’ inside the ‘greater than’ condition. This is incorrect as the runoff should be rainfall minus absorption limit, not the other way around…
In the example, the LLM identifies the pertinent text in the student’s response, discerns that the student is incorrect, explains why (using the rubric), and awards the correct score (0). This type of feedback is informative to teachers and students, explaining why a score was awarded instead of merely returning a numerical value. Providing scoring transparency instills more trust in the LLM and automated essay scoring systems if stakeholders understand the rationale behind the LLM’s decisions [41]. However, while the LLM’s feedback was generally accurate and informative, our experiments show it could be unpredictable and prone to illogical feedback:
Student: If rainfall is equal to Absorption limit, set absorption to absorption limit. Set runoff
GPT-4: …This is incorrect as the student should set absorption equal to either rainfall or absorption limit in the ‘equal to’ condition per the rubric’s guidance…
As the example shows, the student correctly identified R4 and set absorption to the absorption limit. Interestingly, the LLM cited the correct portion of the student’s response and acknowledged the rubric’s directive but did not award a point. This is a clear case of logical inconsistency (i.e., LLM outputs that exhibit logical contradictions or are inconsistent with their CoT reasoning chains [42]), as the LLM disregards acknowledged information when scoring. This demonstrates the LLM can produce inaccurate scores even when the correct score is obvious to a human. While rare, this also highlights that the LLM can arrive at the wrong answer even when citing the correct portion of the student’s response.
Utilizing CoTAL for the Rules Task demonstrated the LLM’s proficiency in generating accurate scoring predictions and delivering valuable feedback when prompted effectively. This indicates that the LLM can serve as a robust tool for scoring and feedback, especially when CoTAL is employed in conjunction with clearly defined formative assessment questions and rubrics.
V-B Debugging Task
The performance comparison between CoTAL and the Baseline for the Debugging Task appears in Table V. For this task, the LLM needed access to the fictional student’s erroneous computational model (see Section III), which we distilled into textual form for inclusion in our prompt. Token limitations in GPT-4’s context window inhibited Active Learning in the Debugging Task, so the results only include the Response Scoring and Prompt Development components of CoTAL Phase II.
Error Baseline CoTAL
D1 0.178 0.404
D2 0.848 0.926
D3 0.374 0.608
D4 0.820 0.914
D5 0.615 0.678
Total Score 0.561 0.779
TABLE V: CoTAL versus the Baseline Performance for the Debugging Task using Cohen’s QWK. Each “D” corresponds to a different subscore for the Debugging Task and is explained in Table II. Total Score compares the LLM’s prediction of the total score (i.e., the sum of all 5 subscores) to the human-assigned score. For each metric, the best-performing implementation is in boldface.
The Debugging Task Baseline implementation resulted in an average QWK of 0.567, while applying CoTAL generated an average QWK of 0.706, representing a 24.5% increase. Baseline performance for the Debugging Task was worse than the Rules Task, with only 2 of 5 subscores having QWKs $\geq 0.8$. This was due to the LLM’s tendency to generate false positives, which outnumbered false negatives 19 to 2. The LLM’s Total Score predictions did not suffer from the “one under” issue but did not align as well with the human scorers (QWK$=0.561$). The LLM correctly predicted the Total Score in 23/33 instances, with all 10 incorrect predictions being overscoring due to false positives. Six of these incorrect predictions were overscored by 2 or more points.
CoTAL produced significantly better results, even without Active Learning. All Debugging Task subscores and Total Scores improved. The false positive to false negative ratio dropped from 19:2 with the Baseline to 2:1 using CoTAL. The severity of the imbalance was reduced, and the number of incorrect predictions decreased, except for D5. The LLM produced five fewer incorrect predictions across subscores with CoTAL than with the Baseline. For Total Score, QWK rose by 0.218 (a 38.9% gain), and the number of instances incorrectly scored by two or more points dropped from 6 to 2, indicating better alignment with human scorers. The LLM provided feedback linked to the student’s response, rubric, prompt guidelines, and computational model using CoTAL. An example of successful LLM feedback for D1 is shown below.
Student: The set absorption limit ( inch ) of the selected material should be right under set rainfall to 1.
GPT-4: …This indicates the student correctly identifies the absorption limit is initially set in the wrong part of the code and that the absorption limit should be set before the first ‘if’ statement…
In the example, the LLM isolates the specific portion of the student’s response that corresponds to D1. The LLM cites the student, explains why the response earns a point according to the rubric and provides the correct score (1). This example highlights the LLM’s ability to identify relevant portions of the student’s response, connect it to the rubric, and assign the correct score. The LLM often referenced other prompt components, such as guidelines and the computational model, citing specific line numbers and code blocks to clarify its scoring decisions. This demonstrates that CoTAL enables the LLM to discern meaning from textual representations of block-based code, even without Active Learning.
Although the LLM effectively interpreted the computational model for scoring and generating feedback, it sometimes struggled with the extra context. As a result, the LLM occasionally had trouble tracking all relevant information. The example below depicts an instance where the LLM incorrectly awarded a point for D3.
Student: In the third rule, it should be rainfall instead of the absorption.
GPT-4: The student says, ‘In the third rule, it should be rainfall instead of the absorption.’ The third rule is the ‘greater than’ condition, which is represented in the code by the following block on Line 8: ‘set [Absorption (inch)] to [Absorption Limit (inch)]’. Therefore, the student correctly identifies that in the ‘less than’ condition, absorption should actually be set to rainfall per the Absorption Set to Absorption Limit rubric criteria…
In this example, the LLM correctly identified the “third rule” (from the Rules Task) as the “greater than” condition but incorrectly used it to award a point for D3, which belongs to the “less than” condition. The LLM linked the “greater than” condition to the “set [Absorption (inch)] to [Absorption Limit (inch)]” code block on Line 8, though the “greater than” condition did not appear until Line 10. It then referenced the rubric for D3 and the “less than” condition before erroneously awarding a point for D3. This is a case of context inconsistency (i.e., LLM outputs that are unfaithful to the user’s provided context where the LLM contradicts a fact in the prompt [42]), where the LLM mixed up “rules,” corresponding “if statements,” and line numbers, demonstrating difficulties in tracking multiple pieces of information. This example also highlights the interconnectedness of our formative assessments, rubrics, and prompts, requiring domain knowledge from the Rules Task to assess and provide feedback for Debugging Task responses.
Despite some context inconsistency errors, the LLM computed the correct score and provided meaningful feedback for most instances using CoTAL. Additionally, CoTAL demonstrated the LLM’s ability to consider textual representations of block-based code. The Debugging Task also identified opportunities for refining rubrics and formative assessment questions. For instance, students and the LLM often conflated absorption and absorption limit in the Rules and Debugging Tasks. This finding can help refine our formative assessment questions, rubrics, and prompts by clarifying this distinction. We hypothesize that instructing students to distinguish between absorption and absorption limit will improve their understanding of science and computing concepts.
V-C Engineering Task
Unlike the previous Rules and Debugging Tasks, the Engineering Task rubric did not have multiple subscores. Instead, students received a single score in the range 0 to 4 (inclusive) based on the rubric in Table III. QWK for the Baseline was 0.970, while QWK using CoTAL was 0.965. The Engineering Task Baseline resulted in four incorrect scoring predictions out of 32, and the QWK score of 0.97 was comparable to the agreement between the two humans during IRR (QWK = 0.975). This supports our previous hypothesis that the human scorer agreement during IRR may indicate the LLM’s initial alignment with human consensus [12].
CoTAL reduced incorrect predictions from 4 to 2, with QWK performance falling slightly by 0.005 due to one instance deviating from the human score by more than one point. This reveals two issues: (1) a decrease in QWK doesn’t imply reduced accuracy, as CoTAL achieved higher accuracy with fewer errors; and (2) prompt engineering may yield diminishing returns for tasks where the LLM performs well out-of-the-box. The performance improvement with CoTAL is attributed to Active Learning, where the LLM initially overlooked students’ discussions on engineering constraints. Students mentioned specific considerations like “cost” and “accessible squares,” but the LLM did not connect these to the rubric. After integrating a validation set instance into the prompt and using corrective CoT reasoning chains, the LLM’s accuracy improved.
Similar to the Rules and Debugging Tasks, the LLM assigned accurate scores. It provided helpful feedback using CoTAL to cite evidence from students’ responses and connected it with the rubric. For example, when a student answered “yes” in the Engineering Task (incorrect answer; 0 points), the LLM still offered a constructive critique:
Student: Yes. Morgan’s second design is better because its cheaper and has less runoff.
GPT-4: …While this indicates the student understands the design constraints (Runoff_Constraint and Cost_Constraint, in this case), the student’s answer is ‘Yes’…the student earned a score of 0.
This example illustrates that the LLM correctly awarded a score of 0, as the student responded “yes” to the assessment question. However, the LLM’s feedback also highlighted the student’s understanding of key design constraints—cost and runoff—even though no points were earned. This recognition is important: it surfaces evidence of conceptual understanding that might otherwise go unnoticed, helping teachers acknowledge student understanding and providing them with affirming, formative feedback despite an incorrect response. Without this level of explanation, educators might overlook the student’s grasp of domain concepts.
We also noted instances of undesirable LLM behavior. Similar to the Rules and Debugging Tasks (as well as the Science Concepts and Reasoning Task from earlier research), the LLM can be deceived by misleading responses. In one case, a student stated, “Morgan needs to check how other amounts of rainfall affect her design,” for which the LLM awarded 3 points. However, to earn 3 points, students needed to identify the difference in rainfall between tests, not the amounts. This issue was noted during Response Scoring and extensively discussed by the research team. Despite emphasizing this in the rubric and prompt guidelines, the LLM still scored this answer incorrectly.
Another issue was that the LLM sometimes included flawed reasoning in its scoring explanations. In one instance, a student responded: “No. Because she one has better cost and worse absorption, and the other has better absorption and worse cost.” The LLM cited the correct portion of the student’s response about engineering constraints, noting it “shows an understanding of the trade-offs between the Engineering Constraints”. Human scorers were clear that this response scored two points based on the rubric. However, the LLM incorrectly stated, “the student does not mention the different rainfall values or the specific Engineering Constraints by name, so the student cannot be awarded 2, 3, or 4 points.” This is an instruction inconsistency hallucination, where the LLM deviates from an explicit instruction provided by the user [42] (the prompt does not require students to identify constraints “by name” for credit). During Active Learning, we used CoT reasoning to show that responses mentioning trade-offs between absorption, runoff, and cost should receive 2 points. Despite recognizing the student’s understanding of the constraints, the LLM failed to award the correct score of 2 points.
Overall, the LLM effectively scored student responses and provided clear rationales using CoTAL. Like the Rules Task, the Engineering Task showed the LLM’s ability to score and give feedback explicitly linked to the rubrics and student comprehension. This study and previous research typically used binary multi-label scoring, but the Engineering Task demonstrated the LLM’s effectiveness with a multi-class (5-way) scoring scheme. Although CoTAL resulted in a slight drop in QWK score and some hallucinations, it allowed the LLM to maintain near-perfect alignment with human scorers and explain scores accurately based on the rubric.
V-D Answering RQ1
Overall, CoTAL generalized well across all tasks and domains, improving average QWK performance by 10.9% and 24.5% for the Rules and Debugging Tasks, respectively. In the Engineering Task, CoTAL predicted two fewer incorrect instances despite a slight QWK drop. Using CoTAL, the LLM agreed with researcher evaluations for 94.7% of the 550 answers across all three formative assessments’ test sets, resulting in errors on 29 of them, where the LLM either fabricated information or otherwise produced outputs that diverged from the researchers’ preferences. Hallucinations were context-dependent, stemming from individual misunderstandings rather than being domain-specific.
Qualitatively, CoTAL produced LLM outputs that accurately explained scores by citing correct evidence from the student’s response and tying it to the rubrics. Although additional mechanisms are needed to reduce LLM hallucinations, CoTAL improved scoring accuracy and enabled LLMs to provide interpretable scores and explanations across multiple domains and assessments. Every formative assessment required its own context-specific prompt (e.g., including domain concepts, assessment questions, and rubrics), yet the same prompt engineering procedure was effective across science, computation, and engineering without methodological adjustments.
VI Analyzing RQ2: Educator and Student Feedback
RQ2 asked: What do teacher and student input reveal about the effectiveness, actionability, and impact of CoTAL’s formative feedback? We answer this question qualitatively by memoing key findings and using constant comparative analysis to analyze educators’ interviews and students’ surveys, respectively.
VI-A Educator Feedback
We conducted semi-structured interviews with two classroom teachers, asking them to reflect on CoTAL’s scoring of several of their students’ Science Concepts and Reasoning responses. Both teachers reported that CoTAL achieved high scoring accuracy. They recognized the utility of LLMs in enhancing teaching efficiency and identifying student learning gaps, thereby guiding subsequent educational interventions. One educator personally undertook the Science Concepts and Reasoning assessment and received a score of 6 out of 9 from the LLM. She noted CoTAL’s effectiveness in pinpointing and explaining her mistakes and detecting her misunderstanding. The other educator emphasized CoTAL’s (and LLMs’, more generally) potential to reduce teacher bias by evaluating students based solely on their answers, not preconceptions about the student:
…as a teacher, sometimes you drop the ball because you’re like, oh, I know, they meant that, even though they didn’t say it…as a teacher, if I’m grading a student’s paper that I don’t know…I can see all of the things they literally say so much more clearly than if it’s a kid I know…that’s one great thing about it being done by AI.
This educator also underscored the value of collaboration between humans and artificial intelligence in education, particularly in teaching language models to recognize situations that necessitate teacher involvement:
…we could train the AI to alert the teacher to that… that’s where [the LLM] could help the teacher quickly go, ‘oh, here’s a place to grow this kid’s knowledge.’
The two teachers proposed refinements to CoTAL, particularly suggesting that the LLM notify the teachers about students who need additional support and provide feedback that recommends subsequent actions to enhance their learning, such as study topics tailored to their knowledge gaps. One teacher outlined three key functions for an enhanced LLM grading system: (1) offering students constructive feedback for reflection, (2) sharing student performance data with educators, and (3) alerting teachers to notable insights in student submissions. This teacher emphasized the importance of the LLM asking thought-provoking questions to students to evaluate and enhance their conceptual understanding of science topic(s):
…that’s where the AI could eventually ask an inquiring question…there’s your next entry into a discussion.
The second teacher expanded on this concept, advocating for the LLM to prompt students to articulate a deep understanding of scientific concepts and interconnections, as opposed to providing surface-level definitions and general overviews of the subject matter:
It’s like, okay, you got the big concept. But the little details that make it richer.
For the first student, you would want to know…well, I mean, they said the three different sizes mean three different quantities, but what are those quantities…
Both teachers also emphasized the significance of acknowledging student achievements and areas for improvement to ensure that formative feedback encompasses recognition and guidance for further learning. Overall, both teachers were receptive to CoTAL and LLM-guided feedback and were optimistic about LLM use in classrooms going forward. They viewed AI systems as “tools” for teachers to improve student feedback and learning, not as replacements for teachers. One teacher stressed the importance of the partnership between teachers and AI, particularly for more routine tasks, as a means of increasing teachers’ productivity:
…[the AI] doing things that make you more productive as the teacher because that’s the kind of stuff that you can do as a teacher, it just is so time intensive, times every kid…so taking some of that legwork out for the teacher…not replacing what the teacher does just, doing the legwork.
The teachers focused on two primary ways CoTAL feedback could be effective and impactful for students and teachers: (1) feedback that encourages critical thinking and fosters a deep understanding of concepts, and (2) feedback that alerts teachers to students’ misunderstandings and identifies opportunities to expand knowledge. Unlike much of the literature on pedagogical agents, where the agent identifies inflection points and performs interventions, both teachers viewed LLMs as the first step in the intervention pipeline, identifying inflection points but allowing teachers to decide on feedback.
However, one teacher noted that it is often impractical for teachers to deliver individual feedback and expressed a desire for CoTAL to provide this feedback in those instances:
…and that’s where this will be powerful, because giving good feedback is not feasible by the sheer amount you have to give, so it doesn’t get given. So finding ways to get more feedback given to kids is where this can come in and be tremendously helpful. Give feedback that…you know, immediate, that you just physically, literally can’t do as a teacher on everything.
Overall, the teachers view the human-AI partnership as a collaboration where AI can: (1) encourage students; (2) alert teachers to students needing assistance; (3) enable teachers to provide more informed, useful feedback; and (4) offer direct feedback when teachers are unavailable. Unlike students, who addressed several of CoTAL’s shortcomings, the teachers focused almost exclusively on the benefits of CoTAL in classroom settings.
VI-B Student Feedback
We conducted a focus group discussion and survey with a class of 23 students, asking whether they agreed with GPT-4’s scores and explanations using CoTAL to score their Science Concepts and Reasoning formative assessments. We also inquired about the usefulness of CoTAL’s feedback and their confidence in AI grading future assignments. The students were enthusiastic participants, providing a range of positive and negative insights into using LLMs as automated graders.
Overall, 61% of the students found CoTAL’s responses helpful, and 65% expressed confidence that an LLM would score future assignments well. Many students specifically noted that the LLM’s scores and explanations were helpful and accurate, feeling that the LLM adequately understood their responses:
It is helpful because it explains why I was correct and it helped me to understand my score
I liked that it explained throughly [sic] what I did wrong and what I did right.
It understood what i said very well.
Several students mentioned they appreciated the LLM’s objectivity, calling its responses honest, helpful, and not biased (“I Ike [sic] how honest it is”; “It was very honest. It was helpful.”; “I liked it was not biased.”). Even in cases where students disagreed with CoTAL’s scoring decisions (i.e., students felt CoTAL underscored their responses), they still expressed an overall openness to LLM grading by answering “yes” to whether or not they trusted AI systems to score future assignments.
The most frequent comment by students was that CoTAL’s describing how it awarded (or did not award) points and explaining their errors was helpful, particularly concerning the LLM citing evidence and tying it back to the rubric:
It was helpful, because i didn’t realize that i wrote rainfall instead of runoff.
Yes, it was helpful, and I know where I should improve.
…the chat does a relatively good job explaining what I did wrong I think I could work with the feedback to help me learn the content better.
However, some students were not as appreciative of CoTAL’s scoring decisions and explanations, and four students could not list a single thing that CoTAL did well. In general, students’ largest complaint was that CoTAL’s feedback lacked sufficient detail to improve incorrect answers:
I think I would learn better if ChatGPT [^5] All formative assessment responses were evaluated using CoTAL with GPT-4, but students often referred to CoTAL as “ChatGPT” due to its ubiquity. answered the question to show what a better response would be like…I wan’t [sic] it to show me what I did wrong and give me a example.
One way to mitigate this issue is to include examples of ideal responses alongside each scoring explanation so the students can compare their responses to the ground truth (as established by their teacher) and understand the differences in relation to the rubric. Students also suggested making the LLM’s tone less harsh and its responses shorter and less repetitive (“Be less harsh.”; “Stop repetition”; “I like how it explains all the things it had for requirements, but its a little long.”) Like their teachers, students expressed a desire for the LLM to acknowledge their achievements in addition to identifying areas of improvement (“Showing what the student did right”; “I think it can do better with showing what you did right.”).
Regarding the effectiveness, actionability, and impact of CoTAL’s formative feedback, students focused on its ability to explain scoring decisions, which helped them identify the misunderstandings that resulted in incorrect responses. Understanding why the LLM awarded its scores allowed students to connect their response shortcomings to the rubric and address learning gaps. This “explainability” fosters student trust in AI systems, as prior work shows students are often reluctant to trust AI-generated grades without understanding the reasoning, hindering their willingness to act on LLM-generated feedback [43].
VI-C Answering RQ2
Teacher interviews and student surveys revealed several insights into CoTAL’s efficacy and utility in classroom settings. Both groups acknowledged the LLM’s scoring accuracy and ability to explain why responses were correct or incorrect, identifying learning opportunities and interventions. They showed a willingness to accept LLMs in educational settings, particularly with human oversight, aligning with previous findings that ChatGPT enjoys a supportive attitude in academia [44, 45]. However, teachers and students noted the need for enhancements. 39% of the students did not find the feedback helpful, and 35% lacked confidence in the LLM’s grading capabilities. Some students wanted the LLM to pinpoint deficiencies and provide more detailed feedback on weaknesses. Teachers emphasized the need for more actionable feedback and “next steps” to address learning gaps and guide a deeper understanding of science topics.
These insights suggest actionable steps to refine CoTAL. We plan to create prompts that highlight correct and incorrect parts of students’ answers. Recognizing correct elements of an answer fosters greater engagement and trust between students and LLMs. To help students grasp their mistakes, we will provide exemplary answers alongside LLM feedback, clarifying why they did not earn full points and which concepts were misunderstood. CoTAL will identify misunderstandings for deeper discussion and inform teachers so that they may support struggling students. We will enhance stakeholder involvement through participatory design during CoTAL Phase II (Response Scoring, Prompt Development, and Active Learning).
VII Discussion and Conclusions
Our work provides several implications for K-12 educational methodology and practice. Traditional automated grading methods, such as supervised fine-tuning, require large datasets and extensive GPU training. In contrast, prompt engineering techniques like WRVRT [33] do not take into account input from teachers or students and lack evaluation across multiple domains. CoTAL presents a significant advantage by aligning LLMs with teachers effectively and adapting responses to reflect the preferences of individual stakeholders in a generalizable way, utilizing only a small set of demonstration instances to incorporate human insight.
This is particularly valuable for subjective tasks such as feedback generation and formative assessment scoring, where teachers differ in their preferences, making a single fine-tuned LLM unlikely to align well with multiple educators. While much work has explored adapting LLMs for personalizing student interactions — typically through content and feedback generation tailored to individual learning styles [46] — CoTAL addresses a notable gap in learning sciences and technologies methodology by personalizing LLMs for educators.
Teacher and student feedback revealed areas for improving stakeholder alignment and underscored the importance of fostering trust in AI systems before expecting real-world adoption. Students often prefer ChatGPT-assisted (i.e., human-in-the-loop) grading relative to fully automated methods [47, 48]. Jiang et al. (2024) [44] found that, while there is “general optimism” about AI’s potential to enhance teaching, concerns persist regarding ethics, discrimination, and regulatory gaps. Lee and Zhai (2024) [49] identified several “inappropriate” ChatGPT use cases where the LLM hallucinated internet material, further adding ChatGPT could be used in ways that “potentially undermine students’ practical inquiry skills over time.” This highlights the need for additional mechanisms that prioritize enhancing LLM reliability, trust, and curricular alignment—not just accuracy—which our findings reinforce.
Hallucination reduction is essential. Huang et al.’s (2023) [42] survey on LLM hallucinations identified several hallucination causes, discussing reasoning failure where “[an LLM] may struggle to produce accurate results if multiple associations exist between questions,” even in instances where the LLM possesses the necessary knowledge. In our case, this manifested through context inconsistency errors due to the LLM conflating the absorption and absorption limit concepts in the Rules and Debugging Tasks. These concepts are inextricably linked and often appear in similar contexts, which challenges differentiation. Just as students struggled to distinguish between these concepts, we hypothesize that the LLM’s training data reflects similar ambiguities, which we will investigate in future work. Hallucination mitigation strategies such as retrieval-augmented generation (RAG) to ground responses in human-curated data and decoding strategies that prioritize adherence to facts and user instructions similarly warrant future investigation.
Our work is not without limitations. While we demonstrated CoTAL’s generalizability across science, computing, and engineering domains within an integrated STEM+C curriculum, further research is needed to evaluate its performance across additional domains and curricula. This study examines CoTAL’s scoring performance in a post hoc setting without assessing its real-time use in classrooms, limiting our ability to determine its impact on student learning. Additionally, human-in-the-loop prompt engineering can be time-consuming, raising concerns about CoTAL’s scalability and how much it reduces teachers’ workload in real-world classroom settings.
Recently, we conducted a follow-up study in four classrooms, each with approximately 26 students, deploying CoTAL via a formative assessment agent. This agent enabled students to discuss their CoTAL-generated formative assessment scores, helping them understand their mistakes and how to improve their performance moving forward. While a comprehensive analysis of students’ learning gains and agent interactions is forthcoming, we observed substantial time savings with CoTAL compared to our earlier efforts involving human scoring, providing students feedback within hours instead of weeks.
In this paper, we introduced a novel approach to formative assessment scoring, Chain-of-Thought Prompting + Active Learning (CoTAL), which integrates ECD principles, human-in-the-loop prompt engineering, and stakeholder-driven refinement of prompts, assessments, and rubrics. We demonstrated CoTAL’s generalizability in scoring and explaining students’ responses across various types of assessment questions and rubrics and multiple domains. We presented a common framework for formative assessment development, evaluation, and refinement that significantly enhances both the scoring performance and explainability of LLMs for formative assessments. This collaboration between researchers, educators, students, and AI offers promising avenues for improving teacher interventions, enhancing learning outcomes, and advancing both instructional methods and curriculum design.